Acolgrove:

Annabelle and Ezra have a dog that eats 3/4 of a cup of food every morning. Annabelle lost the 1/4 cup so she used the 1/3cup and gave him2/3 of a cup of food. Ezra says that Annabelle did not feed the dog enough food. Annabelle says she gave him more than enough. Who is right? Explain how they can compare the fractions to figure out who is correct.

Acolgrove:

@Aqual

Aqual:

so first we need to figure out the ppercentage of a cup 1/4 is

Aqual:

i mean 1/4?

Acolgrove:

sorry i was offline for a minute

Acolgrove:

okay

Acolgrove:

um the pecnetage would be1 of a 4

Aqual:

.25

Aqual:

what is 2/3?

mxddi3:

well, yes, the way Aqual is showing you is correct however, a quicker way to find this out would be to give the fractions a common denominator. since we have 3 and 4, we will find the LCM, which in this case is 12. So we do: \[\frac{ 1 }{ 4 }= \frac{ ? }{ 12 }\] and \[\frac{ 2 }{ 3 }= \frac{ ? }{ 12 }\]

mxddi3:

from here, you change the fractions. If the one that was originally 1/4 is greater as x/12, than she did not feed the dog extra. make sense?